OpenStudy (nyasia):

A family has two cars. The first car has a fuel efficiency of miles per gallon of gas and the second has a fuel efficiency of miles per gallon of gas. During one particular week, the two cars went a combined total of miles, for a total gas consumption of gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

wait do you have any numbers such as how much mpg each car has and what's the total of miles?

OpenStudy (nyasia):

omg im sorry i copied the question wrong. Here it is: A family has two cars. The first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiency of 25 miles per gallon of gas. During one particular week, the two cars went a combined total of 1775 miles, for a total gas consumption of 50 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (nurali):

1st car DATA: rate = 40 mi/gal ; # of gallons = x gal ; distance = 40x miles 2nd car DATA: 25 mi/gal ; # of gallons = 50-x gal ; distance = 25(50-x) miles Equation: 40x + 25*50-25x = 1775 15x = -1250+1775 x = 35 gallons (gas consumed by 1st car) 50-x = 15 gallons (gas consumed by 2nd car)